OpenStudy (anonymous):

Write 32x5y5 with only one exponent. Use parentheses

OpenStudy (kropot72):

Hint:\[ax ^{b}y ^{b}=a(xy)^{b}\]

OpenStudy (anonymous):

ok so 32(xy)^10

OpenStudy (kropot72):

Not really. You multiplied the exponent (5) by two when it was not required to do so.

OpenStudy (anonymous):

hmmm 32 = 2^?

OpenStudy (anonymous):

so would u multiplie 5*5 which is 25

OpenStudy (kropot72):

No. If you read the hint again you will see that the original exponent stays the same value.

OpenStudy (anonymous):

ok im sorry im really bad at math.so if the exponet stays the same then would it be 35(xy)^5

OpenStudy (kropot72):

Yes, you got it. Also as RolyPoly pointed out 32 = 2^5. So you can replace 32 with the number 2 inside the brackets.

OpenStudy (anonymous):

yes thank you very much

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

So did you get the following: \[(2xy)^{5}\]

OpenStudy (anonymous):

would that be 2*2*2*2*2=32 and would still come out to be 32(xy)^5

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

\[(2xy)^{5}\] is the most concise answer. However you could put 32 outside the brackets instead of putting 2 inside the brackets if you like.

OpenStudy (anonymous):

\[[ (-1)^5 \(-2)^-3]^2\]

OpenStudy (anonymous):

i have to Simplify the expression

OpenStudy (anonymous):

-1^5/-2^-6

OpenStudy (anonymous):

(a^x)^y=a^xy

OpenStudy (anonymous):

so how would i get the answer